# Flux-Activated Resonant Control of a Bosonic Quantum Memory

### v1

@@ I.3 Readout and Purcell filter
Another advantage of the compactness of CPW architecture, realised with the hybrid chip design, is simplicity in integration the readout resonator and the associated the bandpass Purcell filter on the same chip. Our Purcell filter, which is shown in Fig. 5, is strongly coupled capacitively to both the readout resonator and the feed line. With this design, our readout resonator, which we place at 7.51 GHz, has coupling $Q\approx$ 700 without introducing any notable limitations on the the transmon $Q$ factor.
@@ II Device Fabrication
As the device used in this study has considerably larger footprint to the ground plane and the flux line with low-pass microwave filter, compared to typical transmon chips used in bosonic cQED devices. For practical purposes, we only EBL for the transmon pattern and photolithography for all other the parts on the chip, as summarised in the table below.
@@ 3 EBL Al Transmon with the SQUID loop
Table 2: Summary of main fabrication steps. A breakdown of the processes used in fabricating respective on-chip features.
An optical microscope image of the SQUID right after EBL exposure is shown in Fig. 8(a).
@@
In the following step, we etch the surface for 60 seconds in Ar ($85\,\%$) – O2 ($15\,\%$) environment and evaporate 115 nm of Nb using Angström Engineering double-angle evaporation system. We perform lift-off soaking the wafer in NMP at 90C for 3 hours and then sonicating in NMP, acetone and methanol 3 minutes each.
@@
Finally, we perform standard electron beam lithography to make the trasmon qubit. The junctions are fabricated using the bridge-free technique, presented in [31]. We then perform double-angle evaporation by first etching the resist at $\pm$20 angles for 15 seconds each. Subsequently, we deposit two layers of Al with thickness of 20 nm and 40 nm Al, separated by oxidation for 20 minutes in 20 mbar O2 environment. We finish the evaporation by capping in 20 mbar O2 environment for 20 minutes.
@@ III.1 Package
Conventional bosonic cQED architecture usually using a monolith package, made of high-purity aluminium, to minimize seam loss [9, 40]. The chip hosting the ancillary elements is inserted into a waveguide to couple capacitively to the cavity mode and the other end of the chip is held by a clamp, which is, in turn, attached to the main package. However, this approach is not suitable for the device we use here. The on-chip ground plane and requirement for wirebonding necessitate a seam placed near the cavity wall.
Figure 9: Design of the superconducting package. (a) Top view of the main body and the bottom clamp. (b) Design of the top clamp with engineered constrictions that reduces the participation of the cavity field near the ground plane of the chip.
Hence, our superconducting package was carefully re-designed to accommodate these requirements. Unlike conventional architecture, our package consists of three separate parts. The main body contains the cavity and the circular waveguide, where the transmon is inserted (Fig. 9(a)). This part is made of 4N6 aluminum to ensure the lifetime of the coaxial $\lambda/4$-cavity.
The waveguide and clamping structures consist of two pieces, the bottom and top clamps. As we expect the high-Q cavity mode field to be weak in the clamps volume, these parts are made of 6061 aluminum alloy to allow more precise machining and mechanical robustness.
Figure 10: Wirebonding. (a) Image of the chip on the bottom clamp after wirebonding. (b) Connection of the on-chip lines to SMA ports. (c) Bond wires and galvanic connection between the on-chip ground plane and the aluminum bottom clamp are also realised with wirebonding.
The top clamp functions as a lid that covers the chip and shields it from the environment. Importantly, it features a 4mm-long extrusion (Fig. 9(b)), which blocks the cavity field from propagating towards SMA ports outside the heavily filtered CPW line.
The bottom clamp features a 0.5 mm-deep pocket, where the chip is located (Fig. 2(a), main text). We use wirebonding to connect the on-chip ground to the surface of the clamp surrounding it (Fig. 10(a),(c)). The package is then mounted on the MXC flange of the dilution refrigerator, effectively connecting on-chip ground plane to the common fridge ground. Similarly, we use wirebonding to galvanically connect the flux and readout line terminals to flat beryllium-copper pins, that are connected to SMA ports (Fig. 10(b)). The drive signals are supplied to the SMA ports through the coaxial lines of the dilution refrigerator.
@@
To avoid deterioration of the cavity lifetime, our design places the seam as far as possible from the cavity. In order to estimate cavity Q-factor due to the seam loss, we integrated the cavity field over a 0.4 mm-wide ribbon along the seam. Even though we don’t have exact data on the loss tangent of the seam, we could use that method to optimize the design based on the relative changes of the cavity Q.
@@ III.2 Assembly
To improve the contact at the seam in our package, we put indium wire in between the cavity and clamping parts. The quality of this joint can be further improved with welding [38] or more advanced surface treatment [27].
Before mounting the device on the MXC flange of the dilution refrigerator, we cover all the seams between the parts with aluminum tape to ensure that the package is light-tight.
@@ IV Calibration of transitions
The Jaynes-Cummings ladder has a rich spectrum of transitions. We calibrate the frequency and amplitude of each transition by preparing the system in one of the adjacent states and then performing spectroscopy and Rabi oscillation experiments, respectively.
To calibrate the $|3-\rangle\leftrightarrow|4+\rangle$ transition, for example, the system is initialized in $|3$-$\rangle$ with the cavity and transmon in resonance. The transmon is then driven by a pulse with variable frequency and subsequently adiabatically detuned from the cavity. Finally, the transmon, now dispersively coupled to the cavity, is used to measure the parity using a standard Ramsey sequence. A change in parity indicates that drive has excited one of the four allowed transitions to states $|2\pm\rangle$ or $|4\pm\rangle$, which appear as four well-resolved features in the spectroscopy diagram. The desired $|3-\rangle\leftrightarrow|4+\rangle$ transition is identified as the one with highest frequency, in accordance to the theory.
After the transition frequency is identified, the amplitudes of excitation pulses are calibrated with power Rabi experiments. A gaussian pulse with standard deviation $\sigma$ and a total duration of $4\sigma$ is applied to the transmon line with variable amplitude. The amplitude corresponding to a $\pi$ pulse is the point in which the parity of the system is changed from odd to even. This experiment allows us to calculate the Rabi frequency $\Omega_{s}$ of the effective drive $\frac{\Omega_{s}}{2}|4+\rangle\langle 3{-}|$. The corresponding frequencies and amplitudes of each JC transition are calibrated using this method and summarized in Tab. 3.
@@
Table 3: Transition calibration parameters. The table shows the frequency of each relevant transition and its respective maximum Rabi frequency $\Omega_{s}$. $\Omega_{s}$ is defined as the coefficient of the $\frac{\Omega_{s}}{2}|\left(N+1\right)\mp\rangle\langle N\pm|$ matrix element when activated through the transmon drive with input signal voltage of $0.4\,\textnormal{V}$. The parameter $\sigma$ is the standard deviation of the Gaussian $\pi$ pulse calibrated for each transition, with total duration of $4\sigma$. The length of the $|5+\rangle\leftrightarrow|6$-$\rangle$ and $|6-\rangle\leftrightarrow|7+\rangle$ transitions were not optimized.
@@
We use this method to calibrate the gaussian excitation pulses used for cavity state preparation in the main text (Tab. 3). The $\sigma$ of each pulse was swept while tracking the parity of the state after a $2\pi$ rotation. The ideal $\sigma$ is that in which the parity of the final state is as close as possible to the parity of the initial state. This corresponds to a compromise between the leakage and decoherence errors that affect short and long pulses, respectively.
V Preparation of arbitrary bosonic states
@@
We used the resonant control protocol to demonstrate the preparation of arbitrary states in the cavity with an analytical control sequence. The sequence consisted of activating the sidebands sequentially using variable-amplitude, Gaussian-shaped pulses with $4\sigma$ duration, where $\sigma\in\left[64,96\right]\,\textnormal{ns}$. This method allows for straightforward preparation of any Fock state and their superpositions in the cavity, as shown in Fig. 11.
@@
Figure 11: Qudit arbitrary state preparation. Wigner plots showing diverse qudit states prepared by sequentially exciting sideband transitions in the Jaynes-Cummings ladder.
@@
Due to a lack of appropriate quantum amplifier at the specific moment of this experiment, we don’t have single-shot readout of the transmon. Thus, the parity signal was calibrated by comparing it to a $1\textnormal{D}$ Wigner measurement of cavity vacuum $|0\rangle$, which is independently measured beforehand. The Wigner plots closely resemble the target states in shape, though with a reduction in contrast. This indicates that the state preparation is mostly limited by the transmon decoherence.
VI Parametric modulation

### v2

@@ I.3 Readout and Purcell filter
Another advantage of the compactness of the CPW architecture, realised with the hybrid chip design, is its simplicity in integration of the readout resonator and the associated bandpass Purcell filter on the same chip. Our Purcell filter, which is shown in Fig. S1, is strongly coupled capacitively to both the readout resonator and the feed line. With this design, our readout resonator, which we place at 7.51 GHz, has coupling quality factor of $Q\approx$ 700 without introducing any notable limitations on the transmon $Q$ factor.
@@ II Device Fabrication
The device used in this study has considerably larger footprint compared to typical transmon chips used in bosonic cQED devices due to the ground plane and the flux line with low-pass microwave filter. For practical purposes, we only use EBL for patterning the transmon, while photolithography was used for all the other parts of the chip, as summarised in the Tab. II.
@@ 3 EBL Al Transmon with the SQUID loop
Table II: Summary of the main fabrication steps. A breakdown of the processes used in fabricating respective on-chip features.
An optical microscope image of the SQUID right after EBL exposure is shown in Fig. S4(a).
@@
To clean leftover resist before the metal deposition, we apply etching for 60 seconds in Ar ($85\,\%$) – O2 ($15\,\%$) environment and evaporate 115 nm of Nb using Angstrom Engineering double-angle evaporation system. We perform lift-off via soaking the wafer in NMP at 90C for 3 hours and then sonicating in NMP, acetone and methanol for 3 minutes each.
@@
Finally, we perform standard electron beam lithography to make the trasmon qubit. The junctions are fabricated using the bridge-free technique, presented in [39]. We then perform double-angle evaporation by first etching the resist at $\pm$20 angles for 15 seconds each. Subsequently, we deposit two layers of Al with thicknesses of 20 nm and 40 nm, separated by oxidation for 20 minutes in 20 mbar O2 environment. We finish the evaporation by capping in 20 mbar O2 environment for 20 minutes.
@@ III.1 Package
Conventional bosonic cQED architecture usually uses a monolith package, made of high-purity aluminium, to minimise seam loss [10, 55]. The chip hosting the ancillary elements is inserted into a waveguide to couple capacitively to the cavity mode and the other end of the chip is held by a clamp, which is, in turn, attached to the main package. However, this approach is not suitable for the device we use here. The on-chip ground plane and requirement for wirebonding necessitate a seam placed near the cavity wall.
Figure S5: Design of the superconducting package. (a) Top view of the main body and the bottom clamp. (b) Design of the top clamp with engineered constrictions.
Hence, our superconducting package was carefully re-designed to accommodate these requirements. Unlike conventional architecture, our package consists of three separate parts. The main body contains the cavity and the circular waveguide, where the transmon is inserted (Fig. S5(a)). This part is made of 4N6 aluminium to ensure the long lifetime of the coaxial $\lambda/4$-cavity.
The clamping structure consists of two pieces, the bottom and top clamps. As we expect the high-Q cavity mode field to be weak in the clamps volume, these parts are made of 6061 aluminium alloy to allow more precise machining and mechanical robustness.
Figure S6: Wirebonding. (a) Image of the chip on the bottom clamp after wirebonding. (b) Connection of the on-chip lines to th SMA ports. (c) Bond wires and galvanic connection between the on-chip ground plane and the aluminium bottom clamp are also realised with wirebonding.
The top clamp functions as a lid that covers the chip and shields it from the environment. Importantly, it features a 4mm-long extrusion (Fig. S5(b)), which blocks the cavity field from propagating towards SMA ports outside the heavily filtered CPW line.
The bottom clamp features a 0.5 mm-deep pocket, where the chip is located (Fig. 2(a), main text). We use wirebonding to connect the on-chip ground to the surface of the clamp surrounding it (Fig. S6(a),(c)). The package is then mounted on the MXC flange of the dilution refrigerator, effectively connecting on-chip ground plane to the common fridge ground. Similarly, we use wirebonding to galvanically connect the flux and readout line terminals to flat beryllium-copper pins, that are connected to SMA ports (Fig. S6(b)). The drive signals are supplied to the SMA ports through the coaxial lines of the dilution refrigerator.
@@
To avoid deterioration of the cavity lifetime, our design places the seam as far as possible from the cavity. In order to estimate cavity Q-factor due to the seam loss, we integrated the cavity field over a 0.4 mm-wide ribbon along the seam. Even though we don’t have exact data on the loss tangent of the seam, we could use that method to optimise the design based on the relative changes of the cavity Q.
@@ III.2 Assembly
To improve the contact at the seam in our package, we put indium wire in between the cavity and the clamping parts. The quality of this joint can be further improved with welding [51] or more advanced surface treatment [33].
Before mounting the device on the MXC flange of the dilution refrigerator, we cover all the seams between the parts with aluminium tape to ensure that the package is light-tight.
@@ IV Calibration of transitions
The Jaynes-Cummings ladder has a rich spectrum of transitions. We calibrate the frequency and amplitude of each transition by preparing the system in one of the adjacent states and then performing spectroscopy and Rabi oscillation experiments.
For example, to calibrate the $|3-\rangle\leftrightarrow|4+\rangle$ transition the system is initialised in $|3$-$\rangle$ with the cavity and transmon on resonance. The transmon is then driven by a pulse with variable frequency and subsequently adiabatically detuned from the cavity. Finally, the transmon, now dispersively coupled to the cavity, is used to measure the parity using a standard Ramsey sequence. A change in parity indicates that drive has excited one of the four allowed transitions to states $|2\pm\rangle$ or $|4\pm\rangle$, which appear as four well-resolved features in the spectroscopy diagram. The desired $|3-\rangle\leftrightarrow|4+\rangle$ transition is identified as the one with highest frequency, in accordance with the theory.
After the transition frequency is identified, the amplitudes of excitation pulses are calibrated with power Rabi experiments. A gaussian pulse with standard deviation $\sigma$ and a total duration of $4\sigma$ is applied to the transmon line with variable amplitude. The amplitude corresponding to a $\pi$ pulse is the point in which the parity of the system is changed from odd to even. This experiment allows us to calculate the Rabi frequency $\Omega_{s}$ of the effective drive $\frac{\Omega_{s}}{2}|4+\rangle\langle 3{-}|$. The corresponding frequencies and amplitudes of each JC transition are calibrated using this method and summarised in Tab. III.
@@
Table III: Calibration of the parameters of the transitions drives. The table shows the frequency of each relevant transition and its respective maximum Rabi frequency $\Omega_{s}$. $\Omega_{s}$ is defined as the coefficient of the $\frac{\Omega_{s}}{2}|\left(N+1\right)\mp\rangle\langle N\pm|$ matrix element when activated through the transmon drive. The parameter $\sigma$ is the standard deviation of the Gaussian $\pi$ pulse calibrated for each transition, with total duration of $4\sigma$. The length of the $|5+\rangle\leftrightarrow|6$-$\rangle$ and $|6-\rangle\leftrightarrow|7+\rangle$ transitions were not optimised.
@@
We use this method to calibrate the gaussian excitation pulses used for cavity state preparation in the main text. The $\sigma$ of each pulse was swept while tracking the parity of the state after a $2\pi$ rotation. The ideal $\sigma$ is such that the parity of the final state is as close as possible to the parity of the initial state. This corresponds to a compromise between the leakage and decoherence errors that affect short and long pulses, respectively.
V Gate time and lifetime comparison
Here, we compare the resonant JC regime that we implemented to the well-known universal control techniques such as selective number-dependent arbitrary phase (SNAP) gate [32, 18], gradient ascent pulse engineering (GRAPE) [17], and echoed conditional displacement (ECD) [13] in the dispersive regime. We will make the comparison in two folds: in terms of the gate duration for generating arbitrary states on the cavity state up to Fock state $|n\rangle$ and in terms of effective lifetime of the system.
First, let us note that the dispersive regime is a special case of the JC regime where the detuning is much larger than the coupling $g$ ($\Delta=\omega_{q}-\omega_{c}\gg g$). In this regime, the main coupling term reads $\hat{H}_{\text{int}}=-\chi\hat{a}^{\dagger}\hat{a}|e\rangle\langle e|$, where $\chi\approx g^{2}/\Delta$. Notably, the dispersive regime can be further subdivided into two classes based on the ratio between $\chi$ and the fastest decoherence rate in the system, $\textnormal{max}(\Gamma_{\phi},\Gamma_{1},\kappa)$, characterising ancilla dephasing $T_{\phi}=1/\Gamma_{\phi}$ and relaxation $T_{1}=1/\Gamma_{1}$, as well as cavity lifetime $T_{1,\textnormal{cav}}=1/\kappa$. In the strong dispersive regime, the bare nonlinearity $\chi$ is considerably larger than $\textnormal{max}(\Gamma_{\phi},\Gamma_{1},\kappa)$, whereas in the weak dispersive regime $\chi\lesssim\textnormal{max}(\Gamma_{\phi},\Gamma_{1},\kappa)$. SNAP and GRAPE operate in the strong dispersive regime, where typically $\chi\sim 1\>\textnormal{MHz}$. For SNAP, by driving the qubit at a certain frequency, an arbitrary phase can be imparted on the cavity Fock state $|n\rangle$ selectively, which essentially sets the gate time as $T\sim 1/\Omega$, where $\Omega$ is the strength of the qubit drive [32]. For generating an arbitrary state within Fock state $|n\rangle$, it would involve applying the SNAP gate $n$ times, which gives $T\sim n/\Omega$ as the total gate time. However, with multi-tone qubit drives, the operation imparting arbitrary phases for all relevant Fock states can be applied simultaneously, making the gate time faster: $T_{\text{SNAP}}\sim 1/\Omega$, which is independent of $n$. And it is important to point out that for SNAP, selectivity demands $\Omega\ll\chi$, and therefore $T_{\text{SNAP}}\sim 1/\Omega\gg 1/\chi$. Another technique, GRAPE, does not require the strict selectivity of SNAP. Also, the qubit and cavity drives can be simultaneously applied while the qubit and cavity are dispersively coupled. This way, GRAPE is more general compared to SNAP, i.e., the solution space for GRAPE contains that of SNAP. This makes the gate time for GRAPE often faster than multi-tone SNAP, which is set by the dispersive coupling strength $T_{\text{GRAPE}}\sim 1/\chi\approx\Delta/g^{2}$ [47].
The technique of ECD, operating in the weak dispersive regime, enhances the nonlinearity in situ via applying strong resonant microwave drive to displace the bosonic mode far from the origin in the phase space. In this displaced frame, the dispersive four-wave-mixing interaction is transformed into three-wave-mixing, resulting in the interaction term resembling the resonant JC Hamiltonian: $\hat{H}_{\textnormal{int}}^{\textnormal{DF}}=\chi(\alpha\hat{a}^{\dagger}+\alpha^{*}\hat{a})\sigma_{z}/2$, where $\alpha$ is the displacement of the cavity in the phase space. Comparing $\hat{H}_{\textnormal{int}}^{\textnormal{DF}}$ to the Eq. (1) of the main text, one can conclude that ECD leverages the effective interaction strength $g_{\textnormal{eff}}=\chi\alpha$. This bosonic enhancement by a factor of $\alpha$, though powerful, is limited by the criticial cavity photon number [5], leading to the upper bound of the enhanced interaction of $g_{\textnormal{eff}}^{\textnormal{max}}=\sqrt{\chi E_{C}/6\hbar}$, where $E_{C}$ is the capacitive energy of the transmon, approximately equal to the transmon anharmonicity. For the typical transmon designs, $E_{C}/2\pi\hbar$ is in the range of $\textnormal{200}-\textnormal{300 MHz}$ in order to suppress the charge noise. In Ref. [13], where $\chi/2\pi=\textnormal{33 kHz}$ and $E_{C}/2\pi\hbar=\textnormal{193 MHz}$, $g_{\textnormal{eff}}^{\textnormal{max}}\approx\textnormal{1 MHz}$, which sets the lower bound for the gate time of one ECD gate to be $t_{\textnormal{ECD}}\geq 1/g_{\textnormal{eff}}^{\textnormal{max}}$. Furthermore, universal control in this regime involves sequential applications of ECDs and qubit rotations, where the complexity of applying an arbitrary gate of dimension $n$ scales with the number of applications $n$.
On the other hand, protocol for generating arbitrary superpositions of Fock states has also been realised utilizing sequential off- and on-resonant (JC regime) qubit-cavity regimes [38, 20]. Here, each JC ladder transition $|e,n-1\rangle\leftrightarrow|g,n\rangle$ has a Rabi rate $\Omega\sqrt{n}$, making each step $t_{n}\sim 1/(g\sqrt{n})$ long. For an operation involving $n$ ladders, the total gate time is approximately $T_{\text{JCS}}=\sum_{n}t_{n}\sim\sqrt{n}/g$. Importantly, this protocol was further improved by incorporating qubit drives with multi-tone frequencies, which drive all relevant transitions at the same time [63], ultimately resulting in $T_{\text{JC}}=\max_{n}t_{n}\sim 1/g$. If we compare this to the dispersive regime (GRAPE), we have $T_{\text{JC}}\sim(g/\Delta)\times T_{\text{GRAPE}}$. Considering that $g/\Delta\approx 0.05-0.2$, the gate time utilizing JC regime is 5205-20 times faster.
For effective lifetime comparison, let us assume that the major contribution comes from qubit energy decay and dephasing, where the corresponding lifetimes are denoted as $T_{1}$ and $T_{\phi}$, since the cavity lifetime is normally much longer. For SNAP, GRAPE, and ECD, the qubit is utilised differently during a particular protocol. However, the error probabilities due to the qubit can be approximated in the same way. In particular, the error probability related to qubit decay is roughly estimated to be $\epsilon_{1}\sim p_{e}\times T/T_{1}$, where $p_{e}$ is the average probability of the qubit being in the excited state (when the decay happens) during the protocol. This approximation also applies for dephasing $\epsilon_{2}\sim 2\sqrt{p_{e}p_{g}}\times T/T_{\phi}$, which is affecting the qubit state in superposition, the maximum of which happens for equal superposition $p_{g}=p_{e}=0.5$, leading to maximum error $\sim T/T_{\phi}$. Therefore, the total error goes as $\epsilon\sim T(p_{e}/T_{1}+2\sqrt{p_{e}p_{g}}/T_{\phi})$. One can define an effective lifetime for the system $T_{\text{sys}}$ in relation to the error and total gate time as $\epsilon\sim T/T_{\text{sys}}$ such that $T_{\text{sys}}\sim(p_{e}/T_{1}+2\sqrt{p_{e}p_{g}}/T_{\phi})^{-1}$.
In the case of the JC regime, as written in the main text the effective decay lifetime of the coupled system is approximately twice that of the qubit $2T_{1}$, while the dephasing (depolarisation) is $T_{\phi}$, as the JC eigenstates contain equal superposition of the qubit’s ground and excited states. Consequently, this results in effective lifetime $T_{\text{sys}}\sim(1/(2T_{1})+1/T_{\phi})^{-1}$ for the system. A comparison for the lifetime is generally hard to make since it is protocol specific. However, one can see that for average qubit probability $p_{e}\sim 0.5$ in the dispersive regime, the system lifetime is equal to the JC regime.
An important point worth stressing is that we detune the qubit after the protocol such that the qubit is decoupled from the cavity, and therefore the cavity’s lifetime is back to its original value. Even though the resonant control protocol introduces the small overhead due to the time required to adiabatically detune the system from resonance, this overhead is constant (does not scale with the gate complexity) and reasonably short even on the timescale of the current system lifetimes ($\sim\textnormal{200}$ ns).
The resonant scheme achieves fast gate time without compromising the lifetime of the system. Therefore, we argue that the JC protocol that we implement here offers the most favourable strategy for universal control of a bosonic mode.
Our method can be further improved by exploring the deep-strong (DS) and ultra-strong (US) coupling regimes of the JC Hamiltonian, providing access to significantly higher values of coupling strength $g$. However, the rotating wave approximation of the quantum Rabi model [49], used in current analysis (see Eq. (1) of the main text), will not be applicable anymore in those regimes. Therefore, new strategies are to be developed for universal control of JC Hamiltonian in DS and US regimes.
Other methods, such as tailored nonlinear Hamiltonians beyond JC dynamics, can also be explored. For instance, the cubic interactions [14] have been demonstrated on a low-Q oscillator by directly terminating the planar cavity with a nonlinear element. The strong nonlinearity inherent to such bosonic mode allows operations on the timescale of tens of nanoseconds. However, such an element is not an effective quantum memory, as the nonlinearity cannot be dynamically decoupled from the cavity, leading to always-on high decoherence rate. Furthermore, this scheme can only operate in the bounded area of the phase space ($|\alpha|<1.5$), outside of which the control is undermined by the static Kerr nonlinearity. Therefore, at this stage such an approach still requires further investigation in order to be a general-purpose tool to achieve universal control of a bosonic mode.
VI Preparation of arbitrary bosonic states
@@
We used the resonant control protocol to demonstrate the preparation of arbitrary states in the cavity with an analytical control sequence. The sequence consisted of activating the sidebands sequentially using variable-amplitude, Gaussian-shaped pulses with $4\sigma$ duration, with the parameters given in the Tab. III. This method allows for straightforward preparation of any Fock state and their superpositions in the cavity, as shown in Fig. S7.
@@
Figure S7: Qudit arbitrary state preparation. Wigner plots showing diverse qudit states prepared by sequentially exciting sideband transitions in the Jaynes-Cummings ladder.
@@
Due to the lack of appropriate quantum amplifier at the specific moment of this experiment, we don’t have single-shot readout of the transmon. Thus, the parity signal was calibrated by comparing it to a $1\textnormal{D}$ Wigner measurement of cavity vacuum $|0\rangle$, which is independently measured beforehand. The Wigner plots closely resemble the target states in shape, though with a reduction in contrast. This indicates that the state preparation is mostly limited by the transmon decoherence.
VII Arbitrary gate on logical states of binomial encoding
As shown in Sec. VI, resonant control protocol enables arbitrary state preparation. Here, we show that an arbitrary gate on the logical states of binomial encoding of the lowest order [50] can be achieved within the framework of the resonant control protocol. This follows as Givens rotations can be utilized to realise $\hat{R}_{x}^{\textnormal{L}}(\theta)$, $\hat{R}_{y}^{\textnormal{L}}(\theta)$, and $\hat{R}_{z}^{\textnormal{L}}(\theta)$ on the subspace $\big\{|\mathord{\downarrow}\rangle^{\textnormal{L}}=|2\rangle,|\mathord{\uparrow}\rangle^{\textnormal{L}}=(|0\rangle+|4\rangle)/\sqrt{2}\big\}$. The rotations can be written explicitly in terms of Givens rotations as
$\hat{R}_{x}^{\textnormal{L}}(\theta)=\hat{\mathcal{R}}_{0,2}(\zeta_{1})\hat{\mathcal{R}}_{2,4}(\zeta_{2})\hat{\mathcal{R}}_{0,2}(\zeta_{1}),$ (S7a)
$\hat{R}_{y}^{\textnormal{L}}(\theta)=\hat{\mathcal{R}}^{-\pi/2}_{0,2}(\zeta_{1})\hat{\mathcal{R}}^{\pi/2}_{2,4}(\zeta_{2})\hat{\mathcal{R}}^{-\pi/2}_{0,2}(\zeta_{1}),$ (S7b)
$\hat{R}_{z}^{\textnormal{L}}(\theta)=\hat{\mathcal{R}}_{0,4}(-2\theta),$ (S7c)
where the parameters $\zeta_{1}=2\arctan\big(\tan(\theta/4)/\sqrt{2}\big)$, $\zeta_{2}=4\arcsin\big(\sin(\theta/4)/\sqrt{2}\big)$, and the notation $\hat{\mathcal{R}}^{\gamma}_{n_{A},n_{B}}(\alpha)$ means Givens rotation between the Fock states $|n_{A}\rangle$ and $|n_{B}\rangle$ by an angle $\alpha$ as described in the main text, where the pulse is phase-shifted by $\gamma$.
Thus, arbitrary control on the logical states of the binomial encoding is enabled via trivial combinations of the Givens rotations.
VIII Parametric modulation